# Likes, Budgets, and Equilibria: Designing Contests for Socially Optimal Advertising

### v1

@@ 3.1 Convergence of Best Response Dynamics
With Assumption 1, we now prove that the BRD converges to a Nash equilibrium, formalized in the following theorem.
@@
Under Assumption 1, the best response dynamics (Algorithm 1) converges to a Nash Equilibrium $\mathbf{b}^{*}$. More precisely, for any given $\epsilon>0$ and any initial point $(\mathbf{b}(0),\boldsymbol{\mu}(0))$ in Algorithm 1, the gradients for $K>\frac{2B^{2}}{\lambda^{2}}\log\left(\frac{\|\mathbf{g}(\mathbf{b}(0),\boldsymbol{\mu}(0)\|_{2}}{\epsilon}\right)$ iterations satisfy $\|(g(\mathbf{b}(K)),\boldsymbol{\mu}(K))\|_{2}<\epsilon$, provided that the step size is selected as $\gamma=\frac{\lambda}{B^{2}}$.
@@ 4 Welfare and Nash Equilibrium
$h_{si}(\mathbf{b})=\frac{e^{kb_{si}}}{\sum_{r\in\mathcal{M}}e^{kb_{ri}}+\delta}$
@@
The NE is given by $\mathbf{b}_{A}^{*}=(0.1496,0.0,0.5263,0.1907,0.1334)$, $\mathbf{b}_{B}^{*}=(0.0337,0.0,0.9663,0.0,0.0)$. The utilities are given by $u_{A}(\mathbf{b}^{*})=1.0431$, $u_{B}(\mathbf{b}^{*})=1.1021$, and therefore $W_{\text{avg}}\left(\mathbf{b}^{*}\right)=1.0431$, $W_{\text{egal}}\left(\mathbf{b}^{*}\right)=1.0726$, $W_{\text{NSW}}\left(\mathbf{b}^{*}\right)=1.0722$.
@@ 4.1 The Optimal Contest Success Function
The CSF $h_{si}(\mathbf{b})=h(b_{si},\mathbf{b}_{-si})$ for each player $s\in\mathcal{M}$ on node $i\in\mathscr{N}$, satisfies the following properties:
@@
    $\frac{\partial^{2}h_{si}}{\partial b_{si}^{2}}<0\quad\forall\,s\in\mathcal{M}.$ (15)
@@
    $\frac{\partial^{2}h_{si}}{\partial b_{si}\partial b_{ri}}<0\quad\forall\,s,r\in\mathcal{M},\ r\neq s.$ (16)
@@
    $\left.\frac{\partial^{2}h_{si}}{\partial b_{si}^{2}}\right|_{\begin{subarray}{c}b_{si}=t+\alpha\\ b_{ri}=t,\,\forall r\neq s\end{subarray}}<\left.\frac{\partial^{2}h_{si}}{\partial b_{si}\partial b_{ri}}\right|_{\begin{subarray}{c}b_{si}=t\\ b_{ri}=t+\alpha\\ b_{ki}=t,\,\forall k\neq r,s\end{subarray}}<0.$ (17)
@@
The above conditions in the assumption constitute the $m$-player extensions of the conditions introduced by Bimpikis et al. (2016) in Proposition 5 to establish the existence and uniqueness of a symmetric Nash equilibrium in their two-player game. Intuitively, the strict concavity condition in (15) ensures that each player’s CSF exhibits diminishing returns with respect to its own investment, implying that the marginal benefit of additional spending decreases as the player allocates more budget to a node. The strategic substitutability condition in (16) captures the negative externalities among competing players, meaning that an increase in one player’s spending reduces the marginal effectiveness of another player’s investment on the same node.
@@
The dominance of diminishing returns condition in (17) contrasts these two opposing effects on firm $s$’s marginal success. The left-hand side captures the diminishing returns effect on a firm $s$ when it increases its own spending from $t$ to $t+\alpha$, while all other m1m-1 firms maintain their spending at $t$. The right-hand side represents the strategic substitutability effect on firm $s$ when a competitor $r$ raises its spending to $t+\alpha$, while firm $s$ and the remaining m2m-2 firms keep their spending fixed at $t$. The inequality thus states that the adverse effect of a firm’s own diminishing returns outweighs the negative strategic impact exerted by any single competitor.
Theorem 2.
Proof.
Bimpikis et al. (2016) (Proposition 5) prove that under assumption 2 (for 2 players), the Nash equilibrium is always unique and symmetric. The proof can be generalized straightforward to $m\geqslant 2$ players as detailed in Appendix˜B.
Since the equilibrium solution is guaranteed to be symmetric always, this means the budget vectors of each firm $s$ are the same, i.e. $b_{si}=b_{i},\;\forall s\in\mathcal{M}$. This implies that the outcome of the contest is same for all the players: $h_{si}(\mathbf{b})=h_{i}(\mathbf{b}),\;\forall s\in\mathcal{M}$, and hence the utilities (3) for all the players are also the same.
$u_{s}(\mathbf{b})=\sum_{i=1}^{n}M_{i}h_{i}(\mathbf{b})-\sum_{i=1}^{n}b_{i}=u(\mathbf{b})$
Since NE is symmetric only, let’s restrict the strategy space to be symmetric profiles only. This reduces the $p$-mean welfare expression (Definition˜2) to be the same as the utility expression:
$W_{p}(\mathbf{b})=\left(\frac{1}{m}\sum_{s=1}^{m}u(\mathbf{b})^{p}\right)^{\frac{1}{p}}=\left(\frac{1}{m}\cdot m\,u(\mathbf{b})^{p}\right)^{\frac{1}{p}}=u(\mathbf{b}).$
The strict concavity (15) of the CSF $h_{si}(\mathbf{b})$ in $b_{si}$ implies that each firm’s utility $u_{s}(\mathbf{b})$ is concave in its own decision vector $\mathbf{b}_{s}$, since it combines a concave total awareness term $\sum_{i=1}^{n}M_{i}h_{si}(\mathbf{b})$ and the negative of a convex (linear) cost term $\sum_{i=1}^{n}b_{si}$. The $p$-mean welfare $W_{p}(\mathbf{b})$, being a concave aggregation of the concave utilities $\{u_{s}(\mathbf{b})\}_{s\in\mathcal{M}}$, is therefore concave in the joint budget profile $\mathbf{b}$, and remains concave in $\mathbf{b}$ when restricted to the symmetric manifold ($b_{si}=b_{i},\ \forall s$).
Since both the planner’s welfare-maximization problem and each firm’s utility-maximization problem are concave programs with convex feasible sets, the KKT conditions are both necessary and sufficient for optimality. Hence, solving the KKT systems yields the optimal welfare-maximizing allocation and the firms’ Nash equilibrium strategies, respectively.
Consider the planner’s constrained maximization game over the symmetric manifold:
$\max_{\mathbf{b}\in\mathbb{R}_{+}^{n},\ \sum_{i}b_{i}\leq C}W_{p}(\mathbf{b}),$
[…]
$\frac{\partial W_{p}}{\partial b_{i}}-\hat{\lambda}+\hat{\mu}_{i}=0,\qquad\hat{\mu}_{i}\geq 0,\;\hat{\lambda}\geq 0,\quad\hat{\mu}_{i}b_{i}=0,\;\hat{\lambda}(C-\sum_{i}b_{i})=0.$
Now consider any single firm $s$ solving its own constrained problem. Its KKT system is identical to the planner’s:
$\frac{\partial u_{s}}{\partial b_{si}}-\lambda_{s}+\mu_{si}=0,\quad\mu_{si}\geq 0,\;\lambda_{s}\geq 0,\;\mu_{si}b_{si}=0,\lambda_{s}(C-\sum_{i}b_{si})=0.$
On symmetric profiles, one has $\partial_{b_{i}}W_{p}(\mathbf{b})=\partial_{b_{si}}u_{s}(\mathbf{b})$ for every $s\in\mathcal{M},i\in\mathscr{N}$, since the derivative of $W_{p}$ equals that of each firm’s payoff evaluated on the symmetric manifold. Hence, setting $\lambda_{s}=\hat{\lambda}$ and $\mu_{si}=\hat{\mu}_{i}$ shows that $(\mathbf{b},\lambda_{s},\mu_{si})$ satisfies each firm’s KKT conditions. Therefore, $\mathbf{b}$ is a best response for every firm, and consequently a symmetric Nash equilibrium of the planner’s welfare-maximizing game.
Corollary 1.
@@
The next proposition illustrates a class of contest success functions that meet the requirements of Assumption 2, through which Theorem˜2 can be implemented.
@@
$h_{si}(\mathbf{b})=\frac{f(b_{si})}{\sum_{s}f(b_{si})+\delta}$ (18)
@@
Now we’ll check if the general concave Tullock function satisfies this simpler implied inequality.
@@
We have shown that for the two-player case, all three conditions of Assumption 2 are satisfied. The proof naturally extends to the $m$-player case as well where $f(x)=f(b_{si})$ and $f(y)$ is replaced by $\sum_{r\neq s}f(b_{ri})$. The strict concavity (15) and the negative mixed partial derivative of the $m$-player CSF (16) follow directly due to similar derivative expressions. For the dominance of diminishing returns condition (17), the key idea is to verify that the inequality holds when comparing any two firms while keeping all others’ expenditures fixed, which can be established analogously to the two-player case. ∎
@@ 5.2 BRD Simulations
We begin by generating a synthetic dataset comprising 100 customers and 10 product holdings. From this dataset, we construct a social network with 100 nodes and estimate its adjacency matrix $\mathscr{E}$, which satisfies our initial assumption of being substochastic. We consider a market with three firms competing over this connected network. The probability of peer-to-peer communication $\alpha$ is set to 0.5. The CSF follows the general functional form given in Proposition 18 (Eq. 18). Specifically, we consider the Tullock (Tullock et al., 1980) CSF $f(b_{si})=b_{si}^{q}$ , which is concave for $q\leqslant 1$ and becomes non-concave for $q>1$. In addition, we analyze two other commonly used functional forms: the concave logarithmic form $f(b_{si})=\log(1+b_{si})$ and the non-concave exponential form $f(b_{si})=\exp(b_{si})$.
@@ 5.3 Results
Figure˜2 presents the evolution of the mean welfare ratio $R(k)$ over BRD iterations for three types of $p$-mean welfare objectives: (i) Nash social welfare ($W_{\text{NSW}}$) , (ii) Average welfare ($W_{\text{avg}}$), and (iii) Egalitarian welfare ($W_{\text{egal}}$). In the concave CSF setting (Figures˜2(a) and 2(c)), as the algorithm converges, $R(k)$ approaches 1 for all three welfare measures, indicating that the equilibrium achieves welfare close to the optimal and hence the $PoA=1$. The shaded regions indicate variability, which is large initially and shrinks as BRD converges, reflecting reduced fluctuations in welfare outcomes over iterations.
In the non-concave CSF setting (Figures˜2(b) and 2(d)), mean welfare ratio $R(k)$ for all three types of social welfare is substantially below the optimum value of $1$. Moreover, the variance across iterations remains high, with fluctuations persisting even in later stages, indicating instability in the best-response dynamics. This inefficiency arises from the lack of concavity in the underlying utility functions, which prevents the BRD algorithm from converging to a stable equilibrium. Consequently, the equilibria obtained are suboptimal, reflecting both poor welfare performance and inconsistent convergence behavior.

### v2

@@ 3.1 Convergence of Best Response Dynamics
These assumptions are the canonical strong-concavity and smoothness conditions used in the analysis of best-response dynamics and gradient-based algorithms (Murhekar et al., 2024). Strong concavity ensures each firm’s utility function is strongly concave in its own strategy, yielding a unique and stable equilibrium, while a bounded norm guarantees that utility gradients do not change abruptly, allowing fixed-step gradient updates to converge.
With Assumption 1, we now prove that the BRD converges to a Nash equilibrium, formalized in the following theorem. In view of Lemma˜1, convergence of BRD to a point where $\|\mathbf{g}(\mathbf{b},\boldsymbol{\mu})\|_{2}$ is arbitrarily small implies convergence to a Nash equilibrium in the limit.
@@
Under Assumption 1, the best response dynamics (Algorithm 1) converges to a Nash Equilibrium $\mathbf{b}^{\star}$. More precisely, for any given $\epsilon>0$ and any initial point $(\mathbf{b}(0),\boldsymbol{\mu}(0))$ in Algorithm 1, the gradients for $K>\frac{2B^{2}}{\lambda^{2}}\log\left(\frac{\|\mathbf{g}(\mathbf{b}(0),\boldsymbol{\mu}(0)\|_{2}}{\epsilon}\right)$ iterations satisfy $\|(\mathbf{g}(\mathbf{b}(K)),\boldsymbol{\mu}(K))\|_{2}<\epsilon$, provided that the step size is selected as $\gamma=\frac{\lambda}{B^{2}}$.
@@ 4 Welfare and Nash Equilibrium
$h_{s,i}(\mathbf{b})=\frac{e^{kb_{s,i}}}{\sum_{r\in\mathcal{M}}e^{kb_{r,i}}+\delta}$
@@
The NE is given by $\mathbf{b}_{A}^{\star}=(0.1496,0.0,0.5263,0.1907,0.1334)$, $\mathbf{b}_{B}^{\star}=(0.0337,0.0,0.9663,0.0,0.0)$. The utilities are given by $u_{A}(\mathbf{b}^{\star})=1.0431$, $u_{B}(\mathbf{b}^{\star})=1.1021$, and therefore $W_{\text{avg}}\left(\mathbf{b}^{\star}\right)=1.0431$, $W_{\text{egal}}\left(\mathbf{b}^{\star}\right)=1.0726$, $W_{\text{NSW}}\left(\mathbf{b}^{\star}\right)=1.0722$.
@@ 4.1 The Optimal Contest Success Function
The CSF $h_{s,i}(\mathbf{b})=h(b_{s,i},\mathbf{b}_{-s,i})$ for each player $s\in\mathcal{M}$ on node $i\in\mathscr{N}$, satisfies the following properties:
@@
    $\frac{\partial^{2}h_{s,i}}{\partial b_{s,i}^{2}}<0\quad\forall\,s\in\mathcal{M}.$ (15)
@@
    $\frac{\partial^{2}h_{s,i}}{\partial b_{s,i}\partial b_{r,i}}<0\quad\forall\,s,r\in\mathcal{M},\ r\neq s.$ (16)
@@
    $\left.\frac{\partial^{2}h_{s,i}}{\partial b_{s,i}^{2}}\right|_{\begin{subarray}{c}b_{s,i}=t+\alpha\\ b_{r,i}=t,\,\forall r\neq s\end{subarray}}<\left.\frac{\partial^{2}h_{s,i}}{\partial b_{s,i}\partial b_{r,i}}\right|_{\begin{subarray}{c}b_{s,i}=t\\ b_{r,i}=t+\alpha\\ b_{ki}=t,\,\forall k\neq r,s\end{subarray}}<0.$ (17)
@@
The above conditions in the assumption constitute the $m$-player extensions of the conditions introduced by Bimpikis et al. (2016) in Proposition 5 to establish the existence and uniqueness of a symmetric Nash equilibrium in their two-player game. Intuitively, the strict concavity condition in Equation˜15 ensures that each player’s CSF exhibits diminishing returns with respect to its own investment, implying that the marginal benefit of additional spending decreases as the player allocates more budget to a node. The strategic substitutability condition in Equation˜16 captures the negative externalities among competing players, meaning that an increase in one player’s spending reduces the marginal effectiveness of another player’s investment on the same node.
@@
The dominance of diminishing returns condition in Equation˜17 contrasts these two opposing effects on firm $s$’s marginal success. The left-hand side captures the diminishing returns effect on a firm $s$ when it increases its own spending from $t$ to $t+\alpha$, while all other m1m-1 firms maintain their spending at $t$. The right-hand side represents the strategic substitutability effect on firm $s$ when a competitor $r$ raises its spending to $t+\alpha$, while firm $s$ and the remaining m2m-2 firms keep their spending fixed at $t$. The inequality thus states that the adverse effect of a firm’s own diminishing returns outweighs the negative strategic impact exerted by any single competitor.
@@
The next proposition illustrates a class of contest success functions that meet the requirements of Assumption 2, demonstrating that such CSFs are feasible design choices.
@@
$h_{s,i}(\mathbf{b})=\frac{f(b_{s,i})}{\sum_{s}f(b_{s,i})+\delta}$ (18)
@@
Now we’ll check if $h(x,y)$ satisfies this simpler implied inequality.
@@
We have shown that for the two-player case, all three conditions of Assumption 2 are satisfied. The proof naturally extends to the $m$-player case as well where $f(x)=f(b_{s,i})$ and $f(y)$ is replaced by $\sum_{r\neq s}f(b_{r,i})$. The strict concavity (15) and the negative mixed partial derivative of the $m$-player CSF (16) follow directly due to similar derivative expressions. For the dominance of diminishing returns condition (17), the key idea is to verify that the inequality holds when comparing any two firms while keeping all others’ expenditures fixed, which can be established analogously to the two-player case. ∎
We now state our main welfare result, establishing that under the proposed CSF design the resulting Nash equilibrium is both well-behaved and socially optimal.
Theorem 2.
Proof.
Bimpikis et al. (2016) (Proposition 5) shows that under Assumption 2, the Nash equilibrium is unique and symmetric for the two-player case. This result extends straightforwardly to $m\geq 2$ players, as detailed in Appendix˜B. Let $\mathbf{b}^{\star}$ denote this unique Nash equilibrium profile.
By symmetry of the equilibrium, each firm allocates its budget identically across nodes, i.e.,
$b^{\star}_{s,i}=b^{\star}_{i},\qquad\forall s\in\mathcal{M}.$
As a result, the contest outcome is identical across firms,
$h_{s,i}(\mathbf{b}^{\star})=h_{i}(\mathbf{b}^{\star}),\qquad\forall s\in\mathcal{M},$
and hence all firms attain the same utility value:
$u_{s}(\mathbf{b}^{\star})=\sum_{i=1}^{n}M_{i}h_{i}(\mathbf{b}^{\star})-\sum_{i=1}^{n}b^{\star}_{i}=:u(\mathbf{b}^{\star}).$
The goal is to characterize Nash equilibria that are also welfare-maximizing. While a welfare-maximizing allocation need not be symmetric in general, only symmetric profiles can coincide with a Nash equilibrium in our setting, since the Nash equilibrium is unique and symmetric. Therefore, we restrict attention to symmetric profiles when comparing Nash equilibrium outcomes with welfare-maximizing allocations.
Under symmetric profiles, the $p$-mean social welfare (Definition˜2) reduces to the common utility value:
$W_{p}(\mathbf{b})=\left(\frac{1}{m}\sum_{s=1}^{m}u_{s}(\mathbf{b})^{p}\right)^{\frac{1}{p}}=\left(\frac{1}{m}\cdot m\,u(\mathbf{b})^{p}\right)^{\frac{1}{p}}=u(\mathbf{b}).$
The strict concavity of the CSF in Assumption 2 implies that each firm’s utility function $u_{s}(\mathbf{b})$ is concave in its own decision vector $\mathbf{b}_{s}$, as it consists of a concave awareness term and a linear cost term. Since the $p$-mean welfare is a concave aggregation of concave utilities, $W_{p}(\mathbf{b})$ is concave in the joint profile $\mathbf{b}$, and remains concave when restricted to the symmetric manifold.
Consider the planner’s welfare-maximization problem restricted to symmetric profiles:
$\max_{\mathbf{b}\in\mathbb{R}_{+}^{n},\ \sum_{i}b_{i}\leq C}W_{p}(\mathbf{b}),$
[…]
$\frac{\partial W_{p}}{\partial b_{i}}(\mathbf{b}^{\mathrm{opt}})-\hat{\lambda}+\hat{\mu}_{i}=0,\qquad\hat{\mu}_{i}\geq 0,\ \hat{\lambda}\geq 0,$
$\hat{\mu}_{i}b^{\mathrm{opt}}_{i}=0,\qquad\hat{\lambda}\Big(C-\sum_{i}b^{\mathrm{opt}}_{i}\Big)=0.$
Now consider any firm $s$ solving its own utility-maximization problem. Its KKT conditions are:
$\frac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b})-\lambda_{s}+\mu_{s,i}=0,\qquad\mu_{s,i}\geq 0,\ \lambda_{s}\geq 0,$
$\mu_{s,i}b_{s,i}=0,\qquad\lambda_{s}\Big(C-\sum_{i}b_{s,i}\Big)=0.$
Evaluating these conditions at the symmetric profile $\mathbf{b}^{\mathrm{opt}}$, we have
$\frac{\partial W_{p}}{\partial b_{i}}(\mathbf{b}^{\mathrm{opt}})=\frac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b}^{\mathrm{opt}}),\qquad\forall s\in\mathcal{M},\ i\in\mathscr{N}.$
Thus, setting $\lambda_{s}=\hat{\lambda}$ and $\mu_{s,i}=\hat{\mu}_{i}$ shows that $\mathbf{b}^{\mathrm{opt}}$ satisfies each firm’s KKT conditions. Hence, $\mathbf{b}^{\mathrm{opt}}$ is a best response for every firm and therefore constitutes a symmetric Nash equilibrium.
Since the Nash equilibrium is unique, we conclude that
$\mathbf{b}^{\star}=\mathbf{b}^{\mathrm{opt}},$
and the equilibrium maximizes the $p$-mean social welfare for all $p\in(-\infty,1]$.
Corollary 1.
Discussions. We conclude by discussing the implications when ˜2 on the CSF is violated. These properties of strict concavity, strategic substitutability, and dominance of diminishing returns, are critical in ensuring that firms’ best responses are well-behaved and that the resulting Nash equilibrium is unique, symmetric, and welfare-maximizing. If the CSF does not satisfy these conditions, the best-response dynamics may fail to converge, or may converge to equilibria that are neither unique nor socially optimal. In particular, the absence of these properties can lead to multiple or asymmetric equilibria (Bimpikis et al., 2016), distorting welfare incentives. We illustrate these failure modes empirically in Section˜5 by simulating alternative CSFs on networks, where deviations from ˜2 lead to degraded convergence and welfare performance.
A key modeling assumption in our framework is that firms have identical total budgets. If firms instead have heterogeneous budgets, the best-response dynamics in Algorithm˜1 remain well-defined and convergence to a Nash equilibrium can still be expected under ˜1, since the update rule and convergence analysis do not depend on the specific value of the budget constraint. However, in this asymmetric setting, the equilibrium need not be unique or symmetric (citation ), and there is no general guarantee that it maximizes social welfare. A complete characterization of equilibrium structure and welfare under heterogeneous budgets would require a separate analysis and is left for future work.
@@ 5.2 BRD Simulations
We begin by generating a synthetic dataset comprising 100 customers and 10 product holdings. From this dataset, we construct a social network with 100 nodes and estimate its adjacency matrix $\mathscr{E}$, which satisfies our initial assumption of being substochastic. We consider a market with three firms competing over this connected network. The probability of peer-to-peer communication $\alpha$ is set to 0.5. The CSF follows the general functional form given in Proposition 18 (Eq. 18). Specifically, we consider the Tullock (Tullock and others, 1980) CSF $f(b_{s,i})=b_{s,i}^{q}$ , which is concave for $q\leqslant 1$ and becomes non-concave for $q>1$. In addition, we analyze two other commonly used functional forms: the concave logarithmic form $f(b_{s,i})=\log(1+b_{s,i})$ and the non-concave exponential form $f(b_{s,i})=\exp(b_{s,i})$.
@@ 5.3 Results
Figure˜1 presents the evolution of the mean welfare ratio $R(k)$ over BRD iterations for three types of $p$-mean welfare objectives: (i) Nash social welfare ($W_{\text{NSW}}$) , (ii) Average welfare ($W_{\text{avg}}$), and (iii) Egalitarian welfare ($W_{\text{egal}}$). In the concave CSF setting (Figures˜1(a) and 1(c)), as the algorithm converges, $R(k)$ approaches 1 for all three welfare measures, indicating that the equilibrium achieves welfare close to the optimal and hence the $PoA=1$. The shaded regions indicate variability, which is large initially and shrinks as BRD converges, reflecting reduced fluctuations in welfare outcomes over iterations.
In the non-concave CSF setting (Figures˜1(b) and 1(d)), mean welfare ratio $R(k)$ for all three types of social welfare is substantially below the optimum value of $1$. Moreover, the variance across iterations remains high, with fluctuations persisting even in later stages, indicating instability in the best-response dynamics. This inefficiency arises from the lack of concavity in the underlying utility functions, which prevents the BRD algorithm from converging to a stable equilibrium. Consequently, the equilibria obtained are suboptimal, reflecting both poor welfare performance and inconsistent convergence behavior.